# Gradient-Based Multi-Objective Deep Learning: Algorithms, Theories, Applications, and Beyond

### v2

@@ 7.1 Reinforcement Learning
Radius Algorithm (RA) [109] employs linear scalarization with fixed preference vectors to transform a MORL problem into a single-objective RL problem. However, a uniform distribution of preferences does not always lead to a uniform distribution of solutions. To address this, various methods focus on improving the configuration of preference vectors. For instance, Optimal Linear Support (OLS) [110, 111] adds new preference vectors dynamically during each iteration. Prediction-guided MORL (PGMORL) [112] employs a hyperbolic model to map preference vectors to objective values, periodically updating preference vectors to optimize both the hypervolume and a diversity indicator.
@@
Although Lu et al. [5] demonstrated that linear scalarization can recover the entire PS when the policy space includes all stochastic policies, this assumption may be too strong in practice, necessitating approaches beyond naive linear scalarization. CAPQL [113] introduces a scalarization method that combines linear scalarization with an entropy term. PPA [114] incorporates a term to align objective values more closely with preference vectors. Several studies [115, 116, 117, 118, 119, 120] have explored other scalarization functions, such as the Tchebycheff function, to identify Pareto-optimal policies. Beyond scalarization, gradient balancing methods (discussed in Section 3.2) also demonstrate promising performance in MORL.
@@
Another research direction focuses on learning the entire Pareto set [121, 122, 123, 124, 125, 126]. In these approaches, the policy network $\pi_{\bm{\theta}}(s,{\bm{\alpha}})$ is conditioned on both the state $s$ and preference vector ${\bm{\alpha}}$. Most of these methods employ training strategies similar to those described in Section 5.2, but with an RL-specific training objective.
7.2 Bayesian Optimization
Bayesian Optimization (BO) [127] is a framework for optimizing expensive-to-evaluate black-box functions by using a probabilistic surrogate model. When extended to handle multiple black-box objectives simultaneously, this approach is known as Multi-Objective Bayesian Optimization (MOBO) [128]. In MOBO, acquisition functions are designed to manage the balance between exploration and exploitation across multiple objectives. Based on this, MOBO methods can be classified into three main categories:
  1. 1.
    Scalarization-based Methods. These methods convert MOO problems into single-objective ones using scalarization functions discussed in Section 2.2 [128, 129]. Zhang and Golovin [130] propose using the hypervolume indicator for scalarization. Instead of scalarizing first and then applying single-objective acquisition functions, methods like SMS-EGO [131], PAL [132, 133], and MOBO-RS [134] extend single-objective acquisitions to multi-objective optimization.
  2. 2.
    Hypervolume Improvement-Based Methods. These methods extend Expected Improvement (EI)  [135] to Expected Hypervolume Improvement (EHVI) [136], which selects candidate points to enhance the current solution set’s hypervolume. Many algorithms are proposed to improve the efficiency and allow parallel candidate generation  [137, 138, 139, 140, 141, 142, 143, 144].
  3. 3.
    Information Theoretic-Based Methods. These methods aim to maximize information gain about the PF by selecting candidate points. PESMO [145] reduces uncertainty in the posterior over the Pareto optimal input set. Some extensions [146, 147, 148, 149] refine this by considering output space. USeMO[150] offers an alternative uncertainty measure based on hyper-rectangle volumes.
@@ 7.2 Bayesian Optimization
The solutions obtained by the aforementioned methods provide discrete approximations of the Pareto set. Lin et al. [151] propose learning a continuous approximation of the Pareto set in MOBO using a hypernetwork (introduced in Section 5.1.1). Subsequent improvements have been proposed to leverage diffusion models [152] and Stein variational gradient descent [153] for efficient solution generation, as well as incorporating data augmentation for Pareto set model building [154].
7.3 Computer Vision
@@
The most representative application of MOO in computer vision is multi-task dense prediction [4], which aims to train a model for simultaneously dealing with multiple dense prediction tasks (such as semantic segmentation, monocular depth estimation, and surface normal estimation) and has been successfully applied in autonomous driving [155, 156]. Since the encoder in computer vision usually contains large number of parameters, sharing the encoder across different tasks can significantly reduce the computational cost but may cause conflicts among tasks, leading to a performance drop in some of the tasks. Hence, many methods are proposed to address the issue from the perspective of MOO, which use scalarization to balance multiple losses [157, 158, 159, 160, 161]. A comprehensive survey on multi-task dense prediction is in [4].
Besides the dense prediction task, MOO is also useful in other computer vision tasks such as point cloud [162, 163, 164], medical image denoising [165], and pose estimation [166].
7.4 Neural Architecture Search
@@
Neural Architecture Search (NAS), which aims to design the architecture of neural networks automatically, has gained significant interest recently [167]. Due to the complex application scenarios in the real world, recent works consider multiple objectives beyond just accuracy. For example, to search an efficient architecture for deployment on platforms with limited resources, many studies incorporate resource-constraint objectives such as the parameter size, FLOPs, and latency. These works are mainly based on gradient-based NAS methods (like DARTS [168]) and formulate it as a multi-objective optimization problem. Among them, [8, 169, 170, 171, 172] employ scalarization to identify a single solution, where the task and efficiency objectives are combined with fixed weights. However, altering the weights of the objectives needs a complete rerun of the search, which is computationally intensive. Therefore, Sukthanker et al. [173] propose to learn a mapping from a preference vector to an architecture using a hypernetwork, enabling to provide the entire PS without the need for a new search.
7.5 Recommendation System
@@
Beyond just focusing on accuracy, recommendation systems often incorporate additional quality metrics such as novelty, diversity, serendipity, popularity, and others [174]. These factors naturally frame the problem as a multi-objective optimization optimization. Traditionally, many methods have employed scalarization techniques to assign weights to the different objectives [175, 176, 177]. In recent developments, gradient-balancing approaches are applied to address multi-objective optimization within recommendation systems to learn a single recommendation model [178, 179, 180, 181]. Additionally, some research efforts aim to identify finite Pareto sets [182] and infinite Pareto sets [183, 184], enabling the provision of personalized recommender models that cater to varying user preferences.
7.6 Large Language Models (LLMs)
@@
Due to the powerful transferability of LLMs, users can fine-tune LLMs to specific downstream tasks or scenarios. This approach separates fine-tuning on each task, causing extensive costs in training and difficulties in deployment. MFTCoder [185] proposes a multi-task fine-tuning framework that enhances the coding capabilities of LLMs by addressing data imbalance, varying difficulty levels, and inconsistent convergence speeds across tasks. CoBa [186] studies multi-task fine-tuning for LLMs and balances task convergence with minimal computational overhead.
Aligning with multi-dimensional human preferences (such as helpfulness, harmfulness, humor, and conciseness) is essential for customizing responses to users’ needs, as users typically have diverse preferences for different aspects. MORLHF [187] and MODPO [188] train an LLM for every preference configuration by linearly combining multiple (implicit) reward models. To avoid retraining, some methods train multiple LLMs separately for each preference dimension and merge their parameters [189, 190] or output logits [191] to deal with different preference requirements. However, these methods need to train and store multiple LLMs, leading to huge computational and storage costs. To improve efficiency, several preference-aware methods are proposed to fine-tune a single LLM for varying preferences by incorporating the relevant coefficients into the input prompts [192, 193, 194] or model parameters [63].
7.7 Miscellaneous
In addition to the applications mentioned above, multi-objective optimization has been applied to a variety of other deep learning scenarios, including meta-learning [195, 196], federated learning [197, 198, 199], long-tailed learning [200, 201, 202], continual learning [203], diffusion models [204, 205, 206, 207], neural combinatorial optimization [208, 122, 209, 210, 211], GFlowNets [212, 213], and physics informed neural networks (PINNs) [214, 215]. These applications highlight the versatility of multi-objective optimization in addressing diverse challenges within deep learning.
8 Resources
In this section, we introduce some benchmark datasets and open-source libraries for gradient-based MOO in deep learning. Benchmark datasets are summarized in Table III, with details in Appendix C.
Two popular libraries, LibMTL [^1] https://github.com/median-research-group/LibMTL [216] and LibMOON [^2] https://github.com/xzhang2523/libmoon [217], provide unified environments for implementing and fairly evaluating MOO methods. Both are built on PyTorch [218] and feature modular designs, enabling flexible development of new methods or application of existing ones to new scenarios. LibMTL [216] mainly focuses on finding a single Pareto-optimal solution and supports $24$ methods introduced in Section 3 and $6$ benchmark datasets. LibMOON [217] mainly focuses on exploring the whole Pareto set. It includes over 20 methods for obtaining a finite set of solutions (introduced in Section 4) or infinite set of solutions (introduced in Section 5).
TABLE III: Summary of benchmark datasets in multi-objective deep learning.
@@ 8 Resources
NYUv2 [219] indoor scene understanding $3$ $1,449$
Taskonomy [220] indoor scene understanding $26$ $\approx 4$M
Cityscapes [221] urban scene understanding $2$ $3,475$
QM9 [222] molecular property prediction $11$ $130$K
Office-31 [223] image classification $3$ $4,110$
Office-Home [224] image classification $4$ $15,500$
CelebA [225] image classification $40$ $\approx 202$K
CIFAR-100 [226] image classification $20$ $60$K
MT10/MT50 [227] reinforcement learning $10/50$
XTREME [228] multilingual learning $9$ $\approx 597$K
PKU-SafeRLHF [229] two-dimensional preference data $2$ $\approx 82$K
UltraFeedBack [230] multi-dimensional preference data $4$ $\approx 64$K
HelpSteer2 [231] multi-attributes preference data $5$ $\approx 21$K
9 Challenges and Future Directions
@@
Theoretical Understanding. While practical methods for MOO in deep learning have seen significant progress, their theoretical foundations remain relatively underexplored.
Most existing research focuses on analyzing the convergence of MOO algorithms to stationary points, while generalization error, which is critical for evaluating performance on unseen data, has received less attention. For instance, Chen et al. [24], as discussed in Section 6.2, provide the algorithm-dependent generalization analysis. Extending this to a broader range of algorithms could offer deeper insights into how different techniques affect generalization.
For Pareto set learning algorithms, there is limited theoretical understanding of how effectively existing algorithms can approximate Pareto sets. Zhang et al. [232] established a generalization bound for HV-based Pareto set learning. However, the effect of various network architectures (Section 5.1) on approximating Pareto sets remain unclear. Additionally, it is uncertain how effective current training strategies (Section 5.2) are in approximating the Pareto set.
@@
10 Conclusion
In this paper, we provide the first comprehensive review of gradient-based MOO in deep learning, covering algorithms, theoretical foundations, applications, benchmark datasets, and software. By unifying diverse approaches and systematically identifying key challenges, this paper serves as a foundational resource to drive future advancements in this important and evolving field.
Appendix A Illustration of Key Concepts in MOO
Figure 6 illustrates the Pareto concepts on the two-objective problem. The blue circles represent Pareto optimal solutions (defined in Definition 2), which indicates that no solution can improve one objective without worsening the other. The blue curve connecting the blue circles denotes the Pareto front, as defined in Definition 3. The yellow circles indicate weak Pareto optimal solutions since they can be improved in one objective without negatively impacting the other. For example, comparing ${\bm{\theta}}_{2}$ and ${\bm{\theta}}_{1}$, ${\bm{\theta}}_{2}$ is a weak Pareto optimal solution since it can be improved in the second objective without affecting the first. The red circles represent solutions that are not Pareto optimal because there exists another solution that strictly Pareto dominates them. For instance, comparing ${\bm{\theta}}_{3}$ and ${\bm{\theta}}_{1}$, ${\bm{\theta}}_{3}$ is not a Pareto optimal solution as ${\bm{\theta}}_{1}$ outperforms it in both objectives.
Figure 6: Illustration of Pareto optimality concepts. The blue, yellow, and red circles denote Pareto optimal solutions, weakly Pareto optimal solutions, and dominated solutions, respectively. The blue curve represents the Pareto front.
Appendix B Additional Details and Further Results for the Experiment in Fig. 4
To demonstrate the differences in the solution distributions produced by various methods, we consider a commonly used benchmark problem, LSMOP1 [84]. Each algorithm is run for 5000 iterations. The solutions obtained by PTML [36], EPO [37, 38], GradHV [47], and MOO-SVGD [48] are shown in Figure 4 in the main paper. Additionally, the solutions generated by MGDA [6], PMGDA [40], UMOD [43], and Smooth Tchebycheff [18] are shown in Figure 7.
From Figures 4 and 7, we can observe that MGDA, which does not incorporate preferences, exhibits poor diversity. The constraint-based approach in PTML produces solutions that are better aligned with the preference vector but still fall short of exact alignment. EPO and PMGDA achieves exact alignment with the preference vector. Some methods that do not use a preference vector, such as GradHV and MOO-SVGD, are also able to produce diverse solutions.
Figure 7: Distributions of solutions obtained by using MGDA [6], PMGDA [40], UMOD [43], and Smooth Tchebycheff [18]. The blue circles denote the solutions, the red curves denote the ground truth PF and the black lines denote the preference vectors.
Appendix C Details of Benchmark Datasets
In this section, we list the details of benchmark datasets in Section 8.
  • NYUv2 dataset [219] is for indoor scene understanding. It has three tasks (i.e., semantic segmentation, depth estimation, and surface normal prediction) with $795$ training and $654$ testing samples.
  • Taskonomy dataset [220] is obtained from 3D scans of about $600$ buildings and contains $26$ tasks (such as semantic segmentation, depth estimation, surface normal prediction, keypoint detection, and edge detection) with about $4$ million samples.
  • Cityscapes dataset [221] is for urban scene understanding. It has two tasks (i.e., semantic segmentation and depth estimation) with $2,975$ training and $500$ testing samples.
  • QM9 dataset [222] is for molecular property prediction with $11$ tasks. Each task performs regression on one property. It contains $130,000$ samples.
  • Office-31 dataset [223] contains $4,110$ images from three domains (tasks): Amazon, DSLR, and Webcam. Each task has 31 classes.
  • Office-Home dataset [224] contains $15,500$ images from four domains (tasks): artistic images, clipart, product images, and real-world images. Each task has $65$ object categories collected under office and home settings.
  • CelebA dataset [225] is a large-scale face attributes dataset, containing $202,599$ face images. Each image has $40$ attribute annotation. Hence, this dataset includes $40$ tasks and each of them is a binary classification problem for one attribute.
  • CIFAR-100 dataset [226] an image classification with $100$ classes, containing $50,000$ training and $10,000$ testing image samples. These $100$ classes are divided into $20$ tasks, each of which is a $5$-class image classification problem.
  • MT10 and MT50 from Meta-World benchmark [227] are two multi-task reinforcement learning settings, containing $10$ and $50$ robot manipulation tasks, respectively.
  • XTREME benchmark [228] aims to evaluate the cross-lingual generalization abilities of multilingual representations. It contains $9$ tasks in $40$ languages, including $2$ classification tasks, $2$ structure prediction tasks, $3$ question-answering tasks, and $2$ sentence retrieval tasks.
  • PKU-SafeRLHF dataset [229] focuses on safety alignment in LLMs, containing $82,118$ question-answering pairs with the annotations of helpfulness and harmlessness.
  • UltraFeedBack dataset [230] contains $63,967$ prompts, each corresponding to four responses from different LLMs. Each response has $4$ aspects of annotations, namely instruction-following, truthfulness, honesty, and helpfulness, generated by GPT-4.
  • HelpSteer2 dataset [231] contains $21,362$ samples. Each sample contains a prompt and a response with $5$ human-annotated attributes (i.e., helpfulness, correctness, coherence, complexity, and verbosity), each ranging between 00 and $4$ where higher means better for each attribute.

### v3

@@ 7.1. Reinforcement Learning
Radius Algorithm (RA) (Parisi et al., 2014) employs linear scalarization with fixed preference vectors to transform a MORL problem into a single-objective RL problem. However, a uniform distribution of preferences does not always lead to a uniform distribution of solutions. To address this, various methods focus on improving the configuration of preference vectors. For instance, Optimal Linear Support (OLS) (Roijers, 2016; Roijers et al., 2014) adds new preference vectors dynamically during each iteration. Prediction-guided MORL (PGMORL) (Xu et al., 2020) employs a hyperbolic model to map preference vectors to objective values, periodically updating preference vectors to optimize both the hypervolume and a diversity indicator.
@@
Although Lu et al. (Lu et al., 2023) demonstrated that linear scalarization can recover the entire PS when the policy space includes all stochastic policies, this assumption may be too strong in practice, necessitating approaches beyond naive linear scalarization. CAPQL (Hu et al., 2023) introduces a scalarization method that combines linear scalarization with an entropy term. PPA (Kyriakis and Deshmukh, 2022) incorporates a term to align objective values more closely with preference vectors. Several studies (Peng and Fain, 2023; Van Moffaert et al., 2013b; Bai et al., 2022; Van Moffaert and Nowé, 2014; Van Moffaert et al., 2013a; Wang, 2024) have explored other scalarization functions, such as the Tchebycheff function, to identify Pareto-optimal policies. Beyond scalarization, gradient balancing methods (discussed in Section 3.2) also demonstrate promising performance in MORL.
@@
Another research direction focuses on learning the entire Pareto set (Pirotta et al., 2015; Lin et al., 2022a; Basaklar et al., 2023; Zhu et al., 2023a; Shu et al., 2024; Liu et al., 2025b). In these approaches, the policy network $\pi_{\bm{\theta}}(s,{\bm{\alpha}})$ is conditioned on both the state $s$ and preference vector $\bm{\alpha}$}. Most of these methods employ training strategies similar to those described in Section 5.2, but with an RL-specific training objective.
7.2. Bayesian Optimization
Bayesian Optimization (BO) (Shahriari et al., 2015) is a framework for optimizing expensive-to-evaluate black-box functions by using a probabilistic surrogate model. When extended to handle multiple black-box objectives simultaneously, this approach is known as Multi-Objective Bayesian Optimization (MOBO) (Knowles, 2006). In MOBO, acquisition functions are designed to manage the balance between exploration and exploitation across multiple objectives. Based on this, MOBO methods can be classified into three main categories: (i) Scalarization-based methods convert MOO problems into single-objective ones using scalarization functions discussed in Section 2.2 (Knowles, 2006; Zhang et al., 2009). Zhang and Golovin (Zhang and Golovin, 2020) propose using the hypervolume indicator for scalarization. Instead of scalarizing first and then applying single-objective acquisition functions, methods like SMS-EGO (Ponweiser et al., 2008), PAL (Zuluaga et al., 2013, 2016), and MOBO-RS (Paria et al., 2020) extend single-objective acquisitions to multi-objective optimization. (ii) Hypervolume improvement-dased methods extend Expected Improvement (EI)  (Jones et al., 1998) to Expected Hypervolume Improvement (EHVI) (Emmerich et al., 2006), which selects candidate points to enhance the current solution set’s hypervolume. Many algorithms are proposed to improve the efficiency and allow parallel candidate generation  (Yang et al., 2019a; Daulton et al., 2020; Yang et al., 2019b; Konakovic Lukovic et al., 2020; Daulton et al., 2021, 2022, 2023; Ament et al., 2023). (iii) Information theoretic-based methods aim to maximize information gain about the PF by selecting candidate points. PESMO (Hernández-Lobato et al., 2016) reduces uncertainty in the posterior over the Pareto optimal input set. Some extensions (Belakaria et al., 2019; Suzuki et al., 2020; Tu et al., 2022; Hvarfner et al., 2022) refine this by considering output space. USeMO(Belakaria et al., 2020) offers an alternative uncertainty measure based on hyper-rectangle volumes.
@@ 7.2. Bayesian Optimization
The solutions obtained by the aforementioned methods provide discrete approximations of the Pareto set. Lin et al. (Lin et al., 2022b) propose learning a continuous approximation of the Pareto set in MOBO using a hypernetwork (introduced in Section 5.1.1). Subsequent improvements have been proposed to leverage diffusion models (Li et al., 2024a) and Stein variational gradient descent (Nguyen et al., 2024) for efficient solution generation, as well as incorporating data augmentation for Pareto set model building (Lu et al., 2024).
7.3. Computer Vision
@@
The most representative application of MOO in computer vision is multi-task dense prediction (Vandenhende et al., 2021), which aims to train a model for simultaneously dealing with multiple dense prediction tasks (such as semantic segmentation, monocular depth estimation, and surface normal estimation) and has been successfully applied in autonomous driving (Ishihara et al., 2021; Liang et al., 2023). Since the encoder in computer vision usually contains large number of parameters, sharing the encoder across different tasks can significantly reduce the computational cost but may cause conflicts among tasks, leading to a performance drop in some of the tasks. Hence, many methods are proposed to address the issue from the perspective of MOO, which use scalarization to balance multiple losses (Xu et al., 2022; Ye and Xu, 2022; Lin et al., 2024a, 2025a; Li et al., 2024b). A comprehensive survey on multi-task dense prediction is in (Vandenhende et al., 2021). Besides the dense prediction task, MOO is also useful in other computer vision tasks such as point cloud (Xie et al., 2023, 2024; Wang et al., 2024b), medical image denoising (Kyung et al., 2024), and pose estimation (Ye et al., 2024c).
7.4. Neural Architecture Search
@@
Neural Architecture Search (NAS), which aims to design the architecture of neural networks automatically, has gained significant interest recently (Ren et al., 2021). Due to the complex application scenarios in the real world, recent works consider multiple objectives beyond just accuracy. For example, to search an efficient architecture for deployment on platforms with limited resources, many studies incorporate resource-constraint objectives such as the parameter size, FLOPs, and latency. These works are mainly based on gradient-based NAS methods (like DARTS (Liu et al., 2019b)) and formulate it as a multi-objective optimization problem. Among them, (Wu et al., 2019; Cai et al., 2019; Wu et al., 2021; Wang et al., 2021a; Yue et al., 2022) employ scalarization to identify a single solution, where the task and efficiency objectives are combined with fixed weights. However, altering the weights of the objectives needs a complete rerun of the search, which is computationally intensive. Therefore, Sukthanker et al. (Sukthanker et al., 2025) propose to learn a mapping from a preference vector to an architecture using a hypernetwork, enabling to provide the entire PS without the need for a new search.
7.5. Recommendation Systems
@@
Beyond just focusing on accuracy, recommendation systems often incorporate additional quality metrics such as novelty, diversity, serendipity, popularity, and others (Zheng and Wang, 2022). These factors naturally frame the problem as a multi-objective optimization optimization. Traditionally, many methods have employed scalarization techniques to assign weights to the different objectives (Di Noia et al., 2017; Patil et al., 2021; Lacerda, 2017). In recent developments, gradient-balancing approaches are applied to address multi-objective optimization within recommendation systems to learn a single recommendation model (Milojkovic et al., 2019; Lin et al., 2019a; Mitrevski et al., 2020; Yang et al., 2024a). Additionally, some research efforts aim to identify finite Pareto sets (Xie et al., 2021) and infinite Pareto sets (Ge et al., 2022; Wilm et al., 2024), enabling the provision of personalized recommender models that cater to varying user preferences.
7.6. Large Language Models (LLMs)
@@
Due to the powerful transferability of LLMs, users can fine-tune LLMs to specific downstream tasks or scenarios. This approach separates fine-tuning on each task, causing extensive costs in training and difficulties in deployment. MFTCoder (Liu et al., 2024a) proposes a multi-task fine-tuning framework that enhances the coding capabilities of LLMs by addressing data imbalance, varying difficulty levels, and inconsistent convergence speeds across tasks. CoBa (Gong et al., 2024) studies multi-task fine-tuning for LLMs and balances task convergence with minimal computational overhead.
Aligning with multi-dimensional human preferences (such as helpfulness, harmfulness, humor, and conciseness) is essential for customizing responses to users’ needs, as users typically have diverse preferences for different aspects. MORLHF (Wu et al., 2023) and MODPO (Zhou et al., 2024a) train an LLM for every preference configuration by linearly combining multiple (implicit) reward models. To avoid retraining, some methods train multiple LLMs separately for each preference dimension and merge their parameters (Rame et al., 2023; Jang et al., 2023) or output logits (Shi et al., 2024) to deal with different preference requirements. However, these methods need to train and store multiple LLMs, leading to huge computational and storage costs. To improve efficiency, several preference-aware methods are proposed to fine-tune a single LLM for varying preferences by incorporating the relevant coefficients into the input prompts (Wang et al., 2024c; Guo et al., 2024; Yang et al., 2024b) or model parameters (Zhong et al., 2024; Lin et al., 2025b).
7.7. Miscellaneous
In addition to the applications mentioned above, multi-objective optimization has been applied to a variety of other deep learning scenarios, including meta-learning (Yu et al., 2023; Wang et al., 2021c), federated learning (Hu et al., 2022; Askin et al., 2024; Kang et al., 2024), long-tailed learning (Li et al., 2024c; Zhao et al., 2024; Zhou et al., 2024b), continual learning (Lai et al., 2025), diffusion models (Hang et al., 2023; Go et al., 2023; Xu et al., 2025a; Yao et al., 2024), neural combinatorial optimization (Li et al., 2020; Lin et al., 2022a; Wang et al., 2023a; Chen et al., 2023c, b), GFlowNets (Jain et al., 2023; Zhu et al., 2023b), and physics informed neural networks (PINNs) (Hwang and Lim, 2024; Liu et al., 2025a). These applications highlight the versatility of multi-objective optimization in addressing diverse challenges within deep learning.
Table 3. Summary of benchmark datasets in multi-objective deep learning.
@@ 7.7. Miscellaneous
NYUv2 (Silberman et al., 2012) indoor scene understanding 33 $1,449$
Taskonomy (Zamir et al., 2018) indoor scene understanding $26$ $\approx 4$M
Cityscapes (Cordts et al., 2016) urban scene understanding 22 $3,475$
QM9 (Ramakrishnan et al., 2014) molecular property prediction $11$ 130130K
Office-31 (Saenko et al., 2010) image classification 33 $4,110$
Office-Home (Venkateswara et al., 2017) image classification 44 $15,500$
CelebA (Liu et al., 2015) image classification $40$ $\approx 202$K
CIFAR-100 (Krizhevsky and Hinton, 2009) image classification 2020 6060K
MT10/MT50 (Yu et al., 2020b) reinforcement learning $10/50$ -
XTREME (Hu et al., 2020) multilingual learning 99 $\approx 597$K
PKU-SafeRLHF (Ji et al., 2024) two-dimensional preference data 22 $\approx 82$K
UltraFeedBack (Cui et al., 2023) multi-dimensional preference data 44 $\approx 64$K
HelpSteer2 (Wang et al., 2024a) multi-attributes preference data 55 $\approx 21$K
8. Resources
In this section, we introduce widely-used benchmark datasets and open-source libraries for gradient-based MOO in deep learning.
8.1. Datasets
Below, we provide details on the benchmark datasets, which are summarized in Table 3:
  • NYUv2 dataset (Silberman et al., 2012) is for indoor scene understanding. It has three tasks (i.e., semantic segmentation, depth estimation, and surface normal prediction) with $795$ training and $654$ testing samples.
  • Taskonomy dataset (Zamir et al., 2018) is obtained from 3D scans of about $600$ buildings and contains $26$ tasks (such as semantic segmentation, depth estimation, surface normal prediction, keypoint detection, and edge detection) with about $4$ million samples.
  • Cityscapes dataset (Cordts et al., 2016) is for urban scene understanding. It has two tasks (i.e., semantic segmentation and depth estimation) with $2,975$ training and $500$ testing samples.
  • QM9 dataset (Ramakrishnan et al., 2014) is for molecular property prediction with $11$ tasks. Each task performs regression on one property. It contains $130,000$ samples.
  • Office-31 dataset (Saenko et al., 2010) contains $4,110$ images from three domains (tasks): Amazon, DSLR, and Webcam. Each task has 31 classes.
  • Office-Home dataset (Venkateswara et al., 2017) contains $15,500$ images from four domains (tasks): artistic images, clipart, product images, and real-world images. Each task has $65$ object categories collected under office and home settings.
  • CelebA (Liu et al., 2015) is a large-scale face attribute dataset comprising $202,599$ face images. Each image is annotated with $40$ binary attributes, resulting in $40$ distinct binary classification tasks.
  • CIFAR-100 (Krizhevsky and Hinton, 2009) is an image classification dataset with $100$ classes, containing 5050K training and 1010K testing images. These $100$ classes are divided into $20$ tasks, each of which is a $5$-class image classification problem.
  • MT10 and MT50 from Meta-World benchmark (Yu et al., 2020b) are two multi-task reinforcement learning settings, containing $10$ and $50$ robot manipulation tasks, respectively.
  • XTREME benchmark (Hu et al., 2020) aims to evaluate the cross-lingual generalization abilities of multilingual representations. It contains $9$ tasks in $40$ languages, including $2$ classification tasks, $2$ structure prediction tasks, $3$ question-answering tasks, and $2$ sentence retrieval tasks.
  • PKU-SafeRLHF dataset (Ji et al., 2024) focuses on safety alignment in LLMs, containing $82,118$ question-answering pairs with the annotations of helpfulness and harmlessness.
  • UltraFeedBack dataset (Cui et al., 2023) contains $63,967$ prompts, each corresponding to four responses from different LLMs. Each response has $4$ aspects of annotations, namely instruction-following, truthfulness, honesty, and helpfulness, generated by GPT-4.
  • HelpSteer2 dataset (Wang et al., 2024a) contains $21,362$ samples. Each sample contains a prompt and a response with $5$ human-annotated attributes (i.e., helpfulness, correctness, coherence, complexity, and verbosity), each ranging between 0 and $4$ where higher means better for each attribute.
8.2. Libraries
Two popular libraries, LibMTL [^1] https://github.com/median-research-group/LibMTL (Lin and Zhang, 2023) and LibMOON [^2] https://github.com/xzhang2523/libmoon (Zhang et al., 2024d), provide unified environments for implementing and fairly evaluating MOO methods. Both are built on PyTorch (Paszke et al., 2019) and feature modular designs, enabling flexible development of new methods or application of existing ones to new scenarios. LibMTL (Lin and Zhang, 2023) mainly focuses on finding a single Pareto-optimal solution and supports $24$ methods introduced in Section 3 and $6$ benchmark datasets. LibMOON (Zhang et al., 2024d) mainly focuses on exploring the whole Pareto set. It includes over 20 methods for obtaining a finite set of solutions (introduced in Section 4) or infinite set of solutions (introduced in Section 5).
9. Challenges and Future Directions
@@
Theoretical Understanding. While practical methods for MOO in deep learning have seen significant progress, their theoretical foundations remain relatively underexplored. Most existing research focuses on analyzing the convergence of MOO algorithms to stationary points, while generalization error, which is critical for evaluating performance on unseen data, has received less attention. For instance, Chen et al. (Chen et al., 2023a), as discussed in Section 6.2, provide the algorithm-dependent generalization analysis. Extending this to a broader range of algorithms could offer deeper insights into how different techniques affect generalization. For Pareto set learning algorithms, there is limited theoretical understanding of how effectively existing algorithms can approximate Pareto sets. Zhang et al. (Zhang et al., 2023a) established a generalization bound for HV-based Pareto set learning. However, the effect of various network architectures (Section 5.1) on approximating Pareto sets remain unclear. Additionally, it is uncertain how effective current training strategies (Section 5.2) are in approximating the Pareto set.
@@
10. Conclusion
In this paper, we provide the first comprehensive review of gradient-based multi-objective deep learning, a field of growing importance as models are required to balance multiple, often conflicting, objectives. We have systematically surveyed the spectrum of algorithms, covering methods for finding a single balanced model, a finite set of models, and an entire infinite Pareto set of models.
The review also delves into the theoretical foundations of these methods, summarizing key results in convergence while highlighting the need for a deeper understanding of generalization. The practical significance of MOO is demonstrated across diverse applications, including its emerging role in Large Language Models (LLMs). By unifying these approaches and identifying key challenges, this paper serves as a foundational resource to guide and inspire future advancements in this critical and rapidly evolving domain.